a charge of 20 x 10-6 c is 20 cm from another charge of 60 x 10-6 c. (b) determine the ep midway between them.


1
Expert's answer
2022-02-25T11:15:48-0500

Answer

Charges are same nature so electric field is in opposite direction.

So net electric field at mid point

"E_p=\frac{9*10^9}{(0.01) ^2}*(60-20) *10^{-6}\\=36*10^6N/C"